# SPVLoc: Semantic Panoramic Viewport Matching for 6D Camera Localization in Unseen Environments

### v1

@@ 4.2 Comparison with the State of the Art
Table 1: Comparison with state-of-the-art baseline on ZInD and S3D: Median translation (terrxy) in cm, rotation error (rerryaw) in deg. for instances localized under 1m. Reporting recall at various translation accuracy levels, recall for inliers ($<$1m and $<$30°), and top-3 recall at 1m. Ours (6D) refers to our model tested with a testset, incorporating pitch and roll rotation variation (±$\pm 10^{\circ}$).
@@
Table 1 shows the results on the datasets ZInD and S3D. Given that we compare against a 2D method, our evaluation also includes recall based on 2D distance and angular error derived from the yaw angle. The yaw angle is extracted from the estimated rotation matrix. Notably, SPVLoc outperforms LASER in terms of localization accuracy, exhibiting higher recall across all categories, an increased inlier rate, and significantly reduced median translation error. In terms of 2D-rotation, our method is slightly less precise. This is because LASER estimates yaw directly while we estimate a full rotation matrix. Estimating camera height and roll and yaw angle is an advantage of our method. In each last column titled 6D, we evaluate our method on an alternate test set with roll and pitch variations (±$\pm 10^{\circ}$). The results suggest that these variations do not substantially impact the accuracy of our method for 2D localization.
Regarding further methods, LASER has showcased a significant performance improvement over PFNet [18] and non-AI-based Monte-Carlo Localization (MCL)[10]. PFNet, originally designed for sequential time updating (such as particle filter), does not perform well in single-query localization scenarios. MCL neglects semantic annotations, uses LIDAR instead of images, which also has been shown to substantially reduce performance [28]. Our substantial improvement over LASER therefore also illustrates a performance gain over aforementioned methods.
@@ 4.3 Performance Studies
Table 2: Ablation Study: SPVLoc variations compared on ZInD with 90-degree FoV. Terr represents xyz, and rerr denotes 3D rotation difference, in contrast to Table 1.
@@
We commence our experimentation with the full network (base) configuration. Subsequently, we modify the network by removing specific components (Table 2). Initially, we test the influence of perspective supervision (-query dec.) and the view segment task head (-view dec.) by excluding them during training. Their exclusion reduces network performance, likely because the ability of the network to understand the image content and establish correspondences is diminished. Removing both (-view/query dec.) further reduces accuracy.
We then exclude the negative samples from different rooms during training (-neg. samples). This significantly reduces the amount of correctly localized images. It supports the hypothesis that teaching the network to recognize spaces where there is no match helps improve the accuracy of predictions.
@@
Figure 5: Qualitative Localization Results: Top to bottom - query, rendering with top-1 estimated pose, panorama with estimated viewport, map. Green box: success for top-1 match. Yellow box: success for top-2 match. Red box: failure case.
Figure 5 displays selected outcomes of SPVLoc. The upper row of the figure showcases the camera images while the images below are rendered from the semantic 3d model with estimated top-1 pose. All green boxed images are correctly localized for the top-1 match. In the yellow box, the top-1 match fails, but the top-2 match is accurate. The red box example cannot be localized. Across all examples, a pose with high visual similarity is found. Ambiguous examples arise from similar layouts occurring in multiple rooms. Failure cases are attributed to either images lacking sufficient structural information for matching or outliers in the data, such as those taken on balconies or in dark cellars. The third row displays the chosen reference panorama, the selected viewport bounding box corresponding to the camera image, and a semi-transparent overlay representing the estimated viewport mask. Additional examples are available in the supplementary material.
@@
Reference positions result from overlaying a grid onto the floorplan, and the grid density significantly affects SPVLoc’s efficiency and inference time. Minimizing references reduces rendering effort and requires the model to estimate larger offsets. E.g. to increase the grid from $0.7m$ to $2m$ reduces the number of samples per $m^{2}$ by a factor of 10.
We vary the grid size to analyze its impact on performance. We also experiment with a local grid per room substituting the global grid (Adv.) reducing the risk of missing rooms entirely. The local grid is centered on the room outline and additionally places one camera in the middle of each opening. Figure 6 shows different grids for an example plan. Table 3 illustrates the relationships among grid size, global and local sampling, and refinement. The optional refinement step, notably enhances the 10cm recall across all configurations. For the largest local grid with refinement, the 1m recall only decreases by 1.18% compared to the baseline configuration. In summary, a trade-off between refinement and dense/sparse scene sampling exists, yet our method consistently delivers strong results with very sparse sampling, making it suitable for large scenes.
@@
The range of pitch and roll angles the network is able to handle is determined by the variations of angles seen during training. We train the network for ZInD using a angle variation of ±$\pm 22^{\circ}$ and test with images where the average variation increases from 00 to $20^{\circ}$ in steps of ±$\pm 5^{\circ}$. Table 5 shows only a slight increase in median errors for rotation and translation as the rotation angles in test images increase. This demonstrates robust estimation, even for larger roll and pitch variations. Assuming a further increase in pitch, there will be a point where images predominantly capture ceiling or floor, making localization impossible.
@@
In large spaces with repeating room layouts, the effectiveness of the approach may be constrained by the limited level of detail in the semantic reference model. Nevertheless, there is an opportunity to enhance the S3D format by incorporating additional semantic classes and more structural information. This could involve the inclusion of features such as staircases, galleries, or even more specific semantic details like sanitary installations. Moreover, input images must at least depict a minimal amount of room structure to be matched.
@@
We split the measured time in scene sampling/preparation time, including the sampling, rendering, and the encoding of the panorama images, and inference/query time, including the encoding of the query image, the feature correlation (including bounding box classification) and the Pose task head. Table 6 shows that our method exhibits a considerably lower query time than LASER [28] and PFNet [18]. This difference places our approach close to real-time for moderately sized scenes. Currently, scene sampling time is slower than [28], mainly due to rendering on the CPU without parallelization [22]. Parallelizing this process using GPU or CPU could significantly accelerate scene preparation.
Table 6: Timing: Performance on ZInD using a NVIDIA Tesla V100, following the configurations outlined in Sec. 2, and compared against results from [28].
@@ 5 Conclusion
Acknowledgements
This work is partially funded by the German Federal Ministry of Economic Affairs and Climate Action (BIMKIT, grant no. 01MK21001H) and the German Federal Ministry for Digital and Transport (EConoM, grant no. 19OI22009C).
@@ Acknowledgements
This supplementary material provides additional insights about our method. Section A includes two additional experiments using the S3D dataset [44]. Section B presents more visual results on ZinD dataset [9] and discussions of specific details of SPVLoc. In the following, literature, figure, and table references refer back to our main paper.
@@ A Additional Experiments
A.1 Refining for a different Aspect Ratio
Table 7: Evaluation of SPVLoc on perspective images from S3D [44]. We compare furnished and unfurnished configuration. We list median errors for 2D localization (translation error terrxy in cm, rotation error rerryaw in degrees) and 6D localization (terrxyz in cm, rerr denoting 3D rotation difference in degrees) for instances localized within 1m. We also report recall at various translation accuracy levels, recall for inliers ($<$1m and $<$30°), and top-3 recall at 1m.
@@ A.1 Refining for a different Aspect Ratio
Perspective 80° FoV   16:9 empty 2D 11.62 1.13 36.55 85.17 87.39 87.08 95.02
6D 12.92 1.49 29.35 85.09 87.34 87.02 95.02
furnished 2D 12.86 1.25 31.68 82.17 85.01 84.83 93.77
6D 14.30 1.61 24.73 82.09 85.01 84.83 93.77
Besides panoramic renderings, S3D [44] also comprises high-resolution perspective images in a 16:9 format with 80 degree horizontal field of view. While we demonstrated a pipeline with perspective training images re-sampled from the 21835 panorama images in the main paper (Table 1), this experiment showcases the usage of the perspective images coming with S3D during training and testing.
Table 7 presents the results for a test set, including 6224 photo-realistically rendered perspective images from S3D test scenes, both with and without furniture. We excluded 181 non-meaningful images that lack any spatial context, showing only a single semantic class, mainly a plain wall. We use our pretrained (Table 1) network and refine it for 10 more epochs with the S3D perspective images. During training, both empty and furnished images were used randomly. The images are zero-padded above and below to fit the square input of the network. With over 85% of the images correctly localized as the top-1 match and over 93% correctly localized as the top-3 match, the network demonstrates its capability to handle the 16:9 test images efficiently. More than 25% of the images are localized correctly within a very small radius of 10 cm.
@@ A.2 Robustness against Furniture
S3D contains renderings of rooms in empty and furnished configuration. Table 7 compares the estimation results for empty and furnished test images, but otherwise equal camera poses. The results indicate, that the trained network is able to efficiently learn to ignore furniture in the images. This can also be seen by the output of the view decoder (see upper branch of Figure  2) in the qualitative results in Figure 7. The decoding of the input image is similar for the empty and furnished image and in both cases similar to the image rendered with ground truth pose. The network suppresses furniture very efficiently. The slight decrease of around 2% in 1-meter recall can mainly be attributed to furniture that obscures the visible room structure and makes matching difficult (Figure 8). The decrease of around 5% in 10-centimeter recall can be explained, for instance, by wardrobes covering room edges, requiring the network to guess the underlying room structure.
@@ B Visual Results
B.1 Field of View Variation
Figure 9 shows an example of SPVLoc correctly estimating the pose of input images with different opening angles, as quantitatively presented in Table 5 in the main paper. Examining the estimated positions on the map reveals that all three images are correctly placed near the ground truth position, which remains consistent across all three instances. Additionally, it is evident that the view rendered from the top-1 camera pose closely matches the input image.
The third column presents a comparison of the viewport estimated with the Segment head of the network (top, white transparent overlay) and the ground truth viewport (bottom), generated via perspective projection. By comparing viewport estimation, viewport ground truth, and the input image, we can observe that the network efficiently learns the perspective projection and establishes visual correspondences between the input and panorama. Importantly, to handle images with varying opening angles, the viewing angle of the camera must be inputted to the Pose head. This implies that knowledge of basic sensor and lens properties is necessary.
B.2 Discussion of Refinement
Table 2 and 3 of the main paper showed how an optional refinement step leads to an improvement in localization recall and pose accuracy. In Figure 10, we show two examples, where the proposed refinement improves the estimated pose. The zoom in the map shows that the estimated pose is closer to the ground truth than the initial estimate. The relative pose offset between selected reference panorama position (pink) and the estimated position (green) is substantially larger than the distance between the estimated position (green) and the refined position (yellow).
As also shown in Table 3, refinement is particularly useful when the distance between reference panorama positions is large, compensating for the cost of rendering an additional panorama during refinement.
During refinement, the query is matched against a single panorama only, leading to quick network inference. Beyond the scope of this paper, our method could also be used for local tracking with a moving camera, where the panorama position is always located at the estimated position of the previous frame.
@@ B.3 Discussion of Viewport Estimation
Figure 11: Viewport score overview: For each example, the top row displays the input image on the left, the model rendered from the estimated top-1 pose in the middle, and a map where each reference position is color-coded relative to the estimated matching score. Here, the ground truth pose is shown with a stripe pattern. For each example, the second row shows three panorama images with high matching scores for the red bounding box and the pixel-based viewport displayed as a semi-transparent white overlay.
To evaluate how well a reference panorama corresponds to an input image, SPVLoc determines a classification score per reference panorama. The method employs anchor-based bounding box estimation [23], where each anchor is classified as either corresponding or not corresponding to the input image. The classification score for a reference panorama is determined as the maximum score over all anchors.
In Figure 11, reference positions are color-coded based on their normalized classification score, ranging between 0 and 1, and displayed in a gradient from red to green. The ground truth camera direction is shown with an arc pattern originating from the ground truth position. The estimated top-3 locations are displayed by smaller transparent circles.
For both examples, we show three reference panoramas with high classification scores. The exact pixel-based viewport estimated by the Segment head is shown as a semi-transparent white overlay.
We see that the reference panoramas achieving high scores capture the content of the input image to a large degree. For example, in Example 1, high scores are achieved for positions where the bay window as well as the wall opening in the direction of the kitchen can be seen in the panorama.
In conclusion, the depthwise-correlation used for the matching demonstrates its ability to incorporate finely-granulated details from the query encoder into the main network branch. It effectively distinguishes subtle distinctions in target images, maps from the photo to the semantic domain, and, in conjunction with the panoramas, successfully induces the correct 6D pose offset. While being based on deep template matching [27], the ability of the operation to bridge the domain gap between semantic and real images as well as dealing with severe perspective distortion is the main novelty of our work. This innovation holds potential for various applications.

### v2

@@ 4.2 Comparison with the State of the Art
Table 1: Comparison with state-of-the-art baseline on ZInD and S3D: Median translation (terrxy) in cm, rotation error (rerryaw) in deg. for instances localized under 1m. Reporting recall at various translation accuracy levels, recall for inliers ($<$1m and $<$30°), and top-3 recall at 1m. Ours refers to our model tested with a slightly different testset, incorporating pitch and roll rotation variation (±$\pm 10^{\circ}$).
@@
Table 1 shows the results on ZInD and S3D. Given that we compare against a 2D method, our evaluation also includes recall based on 2D distance and angular error derived from the yaw angle. The yaw angle is extracted from the estimated rotation matrix. Notably, SPVLoc without refinement outperforms LASER in terms of localization accuracy, exhibiting higher recall across all categories, an increased inlier rate, and significantly reduced median translation error. In terms of 2D rotation, our method is slightly less precise. This is because LASER estimates yaw directly while we estimate a full rotation matrix. Estimating camera height and roll and pitch angle is an advantage of our method. In each last row (Ours), we evaluate our method on an alternate test set with roll and pitch variations (±$\pm 10^{\circ}$). The results suggest that these variations do not substantially impact the accuracy of our method for 2D localization.
Regarding further methods, LASER has showcased a significant performance improvement over PF-net [19] and non-AI-based Monte-Carlo Localization (MCL)[11]. PF-net, originally designed for sequential time updating, does not perform well in single-query localization scenarios. MCL neglects semantic annotations, uses LIDAR instead of images, which also has been shown to drastically reduce performance [30]. Our substantial improvement over LASER therefore also illustrates a performance gain over aforementioned methods. Another experiment using the pre-rendered perspective images of S3D, including a comparison to F3Loc [9], is part of the supplementary material.
@@ 4.3 Performance Studies
Table 2: Ablation study. SPVLoc variations compared on ZInD with $90^{\circ}$ FoV. In contrast to Tab. 1, terr denotes 3D position errors, and rerr denotes 3D rotation errors.
@@
We commence our experimentation with the full network (base) configuration. Subsequently, we modify the network by removing specific components (Tab. 2). Initially, we test the influence of perspective supervision (-query dec.) and the view segment task head (-view dec.) by excluding them during training. Their exclusion reduces network performance, likely because the network’s ability to understand the image content and establish correspondences is diminished. Removing both (-view/query dec.) further reduces accuracy.
We then exclude the negative samples from different rooms during training (-neg. samples). This significantly reduces the amount of correctly localized images. It supports the hypothesis that teaching the network to recognize spaces where there is no match helps to improve the accuracy of predictions.
@@
Figure 5: Qualitative localization results. Top to bottom - query, rendering with top-1 estimated pose, panorama with estimated viewport, map. Green box: success for top-1 match. Yellow box: success for top-2 match. Red box: failure case.
Figure 5 shows selected SPVLoc outcomes. The top row displays camera images, and the second row shows renderings from the semantic 3D model with estimated top-1 poses. All green boxed images are correctly localized for the top-1 match. In the yellow box, the top-1 match fails, but the top-2 match is accurate. The red box example cannot be localized. Across all examples, a pose with high visual similarity is found. Ambiguous examples arise from similar layouts occurring in multiple rooms. Failure cases are attributed to either images lacking sufficient structural information for matching or outliers in the data, such as those taken on balconies or in dark cellars. The third row displays the chosen reference panorama, the selected viewport bounding box corresponding to the camera image, and a semi-transparent overlay representing the estimated viewport mask. Additional detailed visualizations are part of the supplementary material.
@@
Reference positions result from overlaying a grid onto the floorplan, and the grid density significantly affects SPVLoc’s efficiency and inference time. Minimizing references reduces rendering effort and requires the model to estimate larger offsets. E.g. to increase the grid from $0.7m$ to $2m$ reduces the number of samples per $m^{2}$ by a factor of 10. We vary the grid size to analyze its impact on performance. We also experiment with a local grid per room substituting the global grid (Adv.) reducing the risk of missing rooms entirely. The local grid is centered on the room outline and additionally places one camera in the middle of each opening. Figure 6 shows different grids for an example plan. Table 3 illustrates the relationships among grid size, global and local sampling, and refinement. Refinement notably improves 10cm recall across configurations. For the largest local grid with refinement, the 1m recall only decreases by 1.18% compared to the baseline configuration. In summary, a trade-off between refinement and dense/sparse scene sampling exists, yet our method consistently delivers strong results with very sparse sampling, making it suitable for large scenes.
@@
The range of pitch and roll angles the network is able to handle is determined by the variations of angles seen during training. We train the network for ZInD using a angle variation of ±$\pm 22^{\circ}$ and test with images where the range of variation increases from $0^{\circ}$ to $20^{\circ}$ in steps of ±$\pm 5^{\circ}$. Table 5 shows only a slight increase in median errors for rotation and translation as the rotation angles in test images increase. The chosen ranges represent a significant coverage and offer substantial improvements over existing methods. In indoor scenes, larger camera tilt often results in images dominated by the ceiling or floor, rendering them unsuitable for semantic matching.
@@
In large spaces with repeating room layouts, the approach’s effectiveness may be constrained by the limited level of detail in the semantic reference model. However, enhancing the S3D format by adding more semantic classes and structural information, such as staircases or sanitary installations, could reduce ambiguities.
@@
We split the measured time into scene sampling/preparation time, including position sampling, panorama rendering and encoding, and inference/query time, including query image encoding, feature correlation with bounding box classification, and the Pose task head, but no refinement. Table 6 shows that our method exhibits a considerably lower query time than LASER [30] and PF-net [19]. This difference places our approach close to real-time for moderately sized scenes. Currently, scene sampling time is slower than [30], mainly due to rendering on the CPU without parallelization [24]. Parallelizing this process using GPU or CPU could significantly accelerate scene preparation.
Table 6: Timing. Performance on ZInD using a NVIDIA Tesla V100, following the base configuration outlined in Tab. 2, and compared against results from [30].
@@ 5 Conclusion
Acknowledgements.
This work is partially funded by the German Federal Ministry of Economic Affairs and Climate Action (BIMKIT, 01MK21001H) and the German Federal Ministry for Digital and Transport (EConoM, 19OI22009C).
@@ Acknowledgements.
This supplementary material provides additional insights about SPVLoc. Section A includes two additional experiments using the S3D dataset [46]. Section B presents more visual results on ZinD dataset [10] and examines specific aspects of SPVLoc. Section C discusses additional implementation details.
@@ A Additional Experiments
A.1 Refinement for a different Aspect Ratio
Table 7: Evaluation of SPVLoc on perspective images from S3D [46]. We compare furnished and empty configuration. We list median errors for 2D localization (translation error terrxy in cm, rotation error rerryaw in degrees) and 6D localization (terrxyz in cm, rerr denoting 3D rotation difference in degrees) for instances localized within 1m. We also report recall at various translation accuracy levels, recall for inliers ($<$1m and $<$30°), and top-3 recall at 1m.
@@ A.1 Refinement for a different Aspect Ratio
Perspective 80° FoV   16:9 empty 2D 11.70 1.14 35.99 84.06 86.73 86.25 94.79
6D 13.09 1.50 28.77 83.97 86.73 86.21 94.79
furnished 2D 12.86 1.26 30.55 81.58 84.50 84.20 93.36
6D 14.30 1.62 23.54 81.48 84.50 84.18 93.36
Besides panoramic renderings, S3D [46] also comprises high-resolution perspective images in a 16:9 format with 80 degree horizontal field of view. While we demonstrated a pipeline with perspective training images re-sampled from the 21835 panorama images in the main paper, this experiment showcases the usage of the perspective images coming with S3D during training and testing.
Table 7 lists the results for the test sets, including 6405 photo-realistically rendered perspective images from S3D test scenes, both with and without furniture. We use our pretrained network, as in Tab. 1 of the main paper, and refine it for 10 more epochs with the S3D perspective images. During training, both empty and furnished images were used randomly. The images are zero-padded above and below to fit the square input of the network. With over 85% (6D, average of empty and furnished) of the images correctly localized within a 1-meter threshold as the top-1 match and over 94% as the top-3 match, the network demonstrates its capability to handle the 16:9 test images efficiently. More than 26% of the images are localized correctly within a 10 cm threshold.
To our knowledge, the only other work that provides localization results for the perspective images of S3D is F3Loc [9]. This method performs 2D localization and relies solely on a 2D floor plan occupancy without semantic information. Table 8 compares our method with the single-frame variant of F3Loc. Although omitting semantic and height information can be advantageous in certain situations, it significantly reduces the recall rate. The table demonstrates how beneficial using this information is for localization, with nearly four times higher 1m & 30° recall and even greater improvements for 50cm and 10cm recall.
Table 8: Comparison to the single-frame variant of F3Loc[9] using furnished perspective images from S3D. We report recall at different accuracy levels for 2D localization. Results for F3Loc are taken from [9].
Method
Uses
semantics
Reference
model
10cm
50cm
1m
1m & 30°
F3Loc[9] no 2D 1.50 14.60 22.40 21.30
Ours yes 3D 30.55 81.58 84.50 84.20
@@ A.2 Robustness against Furniture
S3D contains renderings of rooms in empty and furnished configuration. Table 7 compares the estimation results for empty and furnished test images, but otherwise equal camera poses. The results indicate, that the trained network is able to efficiently learn to ignore furniture in the images. This can also be seen by the output of the view decoder (the upper branch of Fig. 2 in the main paper) in Fig. 7. The decoding of the input image is similar for the empty and furnished image and in both cases similar to the image rendered with the estimated pose. The network suppresses furniture very efficiently. The slight decrease of around 2% in 1m recall can mainly be attributed to furniture that obscures the visible room structure and makes matching difficult (Fig. 8). The decrease of around 5% in 10cm recall can be explained, for instance, by wardrobes covering room edges, requiring the network to guess the underlying room structure.
@@ B Visual Results
B.1 Discussion of Refinement
The ablation study in the main paper showed how an optional refinement step leads to an improvement in localization recall and pose accuracy. In Fig. 9, we show two examples, where the proposed refinement improves the estimated pose. The zoom in the map shows that the estimated pose is closer to the ground truth than the initial estimate. The relative pose offset between selected reference panorama position (pink) and the estimated position (green) is substantially larger than the distance between the estimated position (green) and the refined position (yellow). As shown in Tab. 3 of the main paper, refinement is particularly useful when the distance between reference panorama positions is large, compensating for the cost of rendering an additional panorama during refinement.
During refinement, the query is matched against a single panorama only, leading to fast inference. Beyond the scope of this paper, our method could also be used for local tracking with a moving camera, where the panorama position is always located at the estimated position of the previous frame.
@@ B.1 Discussion of Refinement
B.2 Field of View Variation
Figure 10 shows an example of SPVLoc correctly estimating the pose of input images with different opening angles, as quantitatively presented in Tab. 5 in the main paper. Examining the estimated positions on the map reveals that all three images are correctly placed near the ground truth position, which remains consistent across all three instances. Additionally, it is evident that the view rendered from the top-1 camera pose closely matches the input image. The third column presents a comparison of the viewport estimated with the Segment head of the network (top, white transparent overlay) and the ground truth viewport (bottom), generated via perspective projection. By comparing viewport estimation, viewport ground truth, and the input image, we can observe that the network efficiently learns the perspective projection and establishes visual correspondences between the input and panorama. Importantly, to handle images with varying opening angles, the viewing angle of the camera must be inputted to the Pose head. This implies that knowledge of basic sensor and lens properties is necessary.
@@ B.3 Discussion of Viewport Estimation
To evaluate how well a reference panorama corresponds to an input image, SPVLoc determines a classification score per reference panorama. The method employs anchor-based bounding box estimation [25], where each anchor is classified as either corresponding or not corresponding to the input image. The classification score for a reference panorama is the maximum score among all anchors.
In Fig. 11, reference positions are color-coded based on their normalized classification score, ranging between 0 and 1, and displayed in a gradient from red to green. The ground truth camera direction is shown with an arc pattern originating from the ground truth position. The estimated top-3 locations are displayed by smaller transparent circles. For both examples, we show three reference panoramas with the highest classification scores. The exact pixel-based viewport estimated by the Segment head is shown as a semi-transparent white overlay. We see that the reference panoramas achieving high scores capture the content of the input image to a large degree. For example, in Example 1, high scores are achieved for positions where the bay window as well as the wall opening in the direction of the kitchen can be seen in the panorama.
In conclusion, the depthwise-correlation used for the matching demonstrates its ability to incorporate finely-granulated details from the query encoder into the main network branch. It effectively distinguishes subtle distinctions in target images, maps from the photo to the semantic domain, and, in conjunction with the panoramas, successfully induces the correct 6D pose offset. While being based on deep template matching [29], the ability of the operation to bridge the domain gap between semantic and real images as well as dealing with severe perspective distortion is one of the main novelties of our work.
@@
Figure 11: Viewport score overview. For each example, the top row displays the input image on the left, the model rendered from the estimated top-1 pose in the middle, and a map where each reference position is color-coded relative to the estimated matching score. Here, the ground truth pose is shown with a stripe pattern. For each example, the second row shows the three panoramas with the highest matching scores, along with the estimated viewport displayed as a semi-transparent white overlay and a red bounding box.
C Further Details
Design of the Matching Module.
Our matching employs feature correlation, where several operations fuse backbone features ($F$ and $R$) from the two branches. Previous studies [6, 29] support the necessity of our module design. The ablation study in [6] shows improved matching by combining element-wise multiplication (there, referred to as cross-correlation) and subtraction with the $N\times 1\times 1$ tensor (justifying $R^{2}$ and $R^{3}$). Building on this, the ablation study in [29] further proves that correlation with the $N\times 3\times 3$ tensor and the concatenation of three tensors enhances matching accuracy (justifying $R^{1}$).
The increase of resolution of $R^{*}$ is attributed to bilinear scaling the features with a small factor. It facilitates linking output features with the pose head by maintaining spatial resolution before flattening.
Negative Samples in Training.
Negative samples are selected across the map, outside of radius $r_{1}$ around target. If the pose is sampled in a different room than the target image, both mask and target scores for all anchors are zero; in the same room, mask and bounding box are calculated as for positive samples, yielding larger offsets. The Pose head is not evaluated for negative samples.
FoV Scaling.
The input image is square, with the assumption that the horizontal and vertical FoV are the same. If an image with a different aspect ratio is used, we recommend zero-padding the centered image to square shape to achieve compatibility with the network (see Sec. A.1). The input for the FoV scaling will then be the larger angle of both. To use the model with zero-padded images, this must be taken into account during training, e.g. by randomly adding symmetrical black bars of different widths at the edges of the image.
Invariance to 3D-Model Inaccuracies.
While ZiND generally annotates ceiling height, it is not always modelled correctly due to inaccuracies and limitations of the data model, e.g., sloping roofs are not modelled in the S3D format, nor are they annotated in ZiND. During training, the network sees images, where mismatches occur and learns to deal with those imperfections e.g. by focusing on other visible elements like windows and doors. If height information is not available at inference time, we recommend using approximated fixed standard heights.
Test Set Design and Comparison with LASER.
We report slightly different results than those of LASER [30] as we conducted experiments using a test set designed to better evaluate pose estimation by requiring contextual information in the images. In contrast to the results of the original LASER experiment, our test images yield more inliers due to their richer semantic variation, albeit with slightly higher median translation errors due to computations over more inliers. LASER’s original experiment involved randomly sampling test yaw angles, resulting in images where the camera faced empty walls, making pose estimation impossible. The authors did not publish the specific angles used for each panorama, whereas we disclose all sampled angles to ensure the reproducibility of our dataset.